OpenStudy (anonymous):

A ball is rolled off the edge of a swimming pool and strikes the water 2.0 m away 0.30 seconds later. What was the initial speed of the ball? its speed when it strikes the water?

OpenStudy (souvik):

at position A the vertical velocity of the ball is zero.. when the ball strikes the water the horizontal velocity does not change as there is no horizontal acceleration.. so \(\huge{v _{x}}\)=2/0.30=20/3 m/s  let the vertical displacement of the ball be \(h\) so \(h\)=1/2\(gt^2\) when the ball strikes the horizontal velocity of the ball be \(v _{y}\) so \(v _{y}=gt\) so the velocity of the ball when it strikes water \(v=\sqrt{v _{x}+v _{y}}\)